OpenStudy (anonymous):

Find the y-intercept for the equation: 3x + 2y = 9

OpenStudy (anonymous):

Substitute x as 0 2y = 9 y = 4.5 the y intercept is therefore (0,4.5)

OpenStudy (anonymous):

y=mx+b b= y intercept 3x + 2y = 9 2y=-3x+9 y=-3/2x+9/2 y intercept: 9/2